# Online Learning–Enhanced Lie Algebraic MPC for Robust Trajectory Tracking of Autonomous Surface Vehicles

### v1

@@ Abstract
Autonomous surface vehicles (ASVs) are easily influenced by environmental disturbances such as wind and waves, making accurate trajectory tracking a persistent challenge in dynamic marine conditions. In this paper, we propose an efficient controller for trajectory tracking of marine vehicles under unknown disturbances by combining a convex error-state MPC on the Lie group with an online learning module to compensate for these disturbances in real time. This design enables adaptive and robust control while maintaining computational efficiency. Extensive evaluations in numerical simulations, the Virtual RobotX (VRX) simulator, and real-world field experiments demonstrate that our method achieves superior tracking accuracy under various disturbance scenarios compared with existing approaches.
@@ I Introduction
The autonomous surface vehicle (ASV) has played an increasingly important role in the application of marine robots, undertaking a wide range of tasks such as hydrographic surveying, environmental monitoring, and water-quality samplings [zereik2018challenges]. For surface vehicles, the time-varying hydrodynamic coefficients and external disturbances pose an obstacle to the controller’s trajectory-tracking ability, especially for underactuated ASV [liu2016unmanned]. For example, winds, waves, currents, and inaccurate dynamics modeling can all disturb accurate tracking trajectory [gao2024dynamic], which can lead to undesired consequences such as grounding, collision risks, or potential structural damage to the hull, thereby degrading reliability and operational safety.
Classical controllers like PID have been applied on marine robots [liu2016unmanned], however, their performance is constrained under strong nonlinearities and time-varying disturbances. In contrast, Model Predictive Control (MPC) has demonstrated effectiveness in trajectory tracking, particularly in more complex environments [cho2020efficient] [veksler2016dynamic]. However, the non-linear hydrodynamics characteristics mean that the computation load of real-time MPC is heavy, resulting in a long response time to control signals, and a low motion frequency [sarda2016station]. Moreover, small computing units lack the capacity to handle such heavy workloads, while large ones consume excessive space and energy, making them impractical for a compact ASV.
Multiple approaches have been put forward to alleviate the computational burden. In [liu2020computationally], MPC with projection neural network performs efficient constrained optimization problems with parallel computational capability. In [annamalai2015robust], adaptive MPC leverages multiple approximated linear models to simplify the calculation. Another approach is to achieve geometric control [ghaffari2022progress], based on a Lie group framework, to exploit symmetry in the problem. In this concept, Lie Algebraic MPC (Lie-MPC) is proposed [teng2022error], and applied on ASV [jang2023convex]. Lie-MPC reduces computation time, making online learning for non-linear hydrodynamics applicable to ASV.
@@
In parallel, several ways of online learning have been proposed. In [tu2019adaptive], a neural network–based adaptive module is incorporated into a model‐based controller to compensate for large wave forces. In [o2022neural],the Deep Neural Network (DNN) is first trained offline to learn the model dynamics, then it is updated online to fit unknown disturbances. Similarly, in [peng2023online], a DNN is used to learn the system dynamics, and an Extended State Observer (ESO) measures error and makes up for residual dynamics estimation, combining them in a new control law.
However, these neural‐network approaches generally require GPU resources for offline pre‐training on large datasets, which increases computational cost. Moreover, like in [peng2023online], the online-learning pipeline is used to directly change the system dynamics, making it difficult to generalize across different system setups. In [zhou2025simultaneous], all disturbances are unified into a single locally Lipschitz term within the MPC dynamics and estimated through feature-based modeling. While this approach has been demonstrated on aerial vehicles and shown to reduce computational burden, its applicability to ASVs and its interaction with a Lie-MPC controller remain unexplored.
In particular, the main contributions of this work are as follows:
@@
The remainder of this paper is organized as follows. Section II provides the preliminaries and mathematical formulations used throughout the paper. Section III presents the details of our proposed method. Experiments are presented in Section IV. Section V concludes the paper and discusses ideas for future studies.
@@ II-A Dynamics on Lie Group
This section briefly introduces the dynamics model on the Lie group $\mathrm{SE}(3)$ and the commonly used notation. For a detailed explanation of Lie groups, refer to [hall2013lie, sola2018micro, bullo2019geometric].
Let $\mathcal{G}$ be an $n$-dimensional matrix Lie group and $\mathfrak{g}$ its associated Lie algebra. For simplicity, we define the following isomorphism
@@
In the specific case of the special Euclidean group $\mathrm{SE}(3)$, the vehicle state can be represented by a rotation matrix $R\in\mathrm{SO}(3)$ and a position vector $p\in\mathbb{R}^{3}$ as
@@
$\xi=\begin{bmatrix}\omega\\ v\end{bmatrix},\quad\xi^{\wedge}=\begin{bmatrix}\omega^{\wedge}&v\\ 0&0\end{bmatrix}\in\mathfrak{se}(3),$ (4)
where $(\cdot)^{\wedge}:\mathbb{R}^{3}\rightarrow\mathfrak{so}(3)$ maps a vector to its corresponding skew-symmetric matrix.
For $X\in\mathrm{SE}(3)$ defined above, the adjoint map has the following matrix representation:
$\mathrm{Ad}_{X}=\begin{bmatrix}R&0\\ p^{\wedge}R&R\end{bmatrix},\qquad\mathrm{ad}_{\xi}=\begin{bmatrix}\omega^{\wedge}&0\\ v^{\wedge}&\omega^{\wedge}\end{bmatrix}$. (5)
The operator $\mathrm{Ad}_{X}$ describes how elements of a Lie group act on its Lie algebra, while $\mathrm{ad}_{\xi}$ represents the Lie bracket operation within the algebra.
@@ II-B Dynamics of ASV in Fossen’s model
We will use a 6-DOF motion equations based on Fossen’s model [fossen1995nonlinear] to describe the ASV, but formulated using Lie group’s ordering of twist vector $\xi=[\omega,v]$. The equations of motion for ASV in the body frame can be described as follows:
$M_{RB}\dot{\xi}$ $+C_{RB}(\xi)\xi$
$+M_{AM}\dot{\xi}_{r}+C_{AM}(\xi_{r})\xi_{r}$
$+D(\xi_{r})\xi_{r}+g(\eta)$
$=\tau_{c}+\tau_{disturb}^{body}$ (6)
where $M_{RB}$ is the rigid-body mass matrix, $C_{RB}(\xi)$ is the rigid-body Coriolis and centripetal matrix, $M_{AM}$ is the added-mass matrix. "Added mass" here represents the inertial hydrodynamic effect, dominated by the pressure field around the hull. It is characterized by a coefficient matrix that relates the ASV’s velocity to the resultant hydrodynamic reaction forces and moments. $C_{AM}(\xi_{r})$ is the added-mass Coriolis and centripetal matrix, $D(\xi_{r})$ is the hydrodynamic damping (drag) matrix. $g(\eta)$ is the hydrostatic restoring force and moment vector, and here we assume the vehicle to be neutrally buoyant, so the hydrostatic term $g(\eta)$ becomes negligible. $\xi$ is the twist vector in the body frame ordered as $\xi=[p,q,r,u,v,w]^{\mathsf{T}}$, $\xi_{r}=\xi-\xi_{c}$ is the relative twist with respect to the ocean current $\xi_{c}$. $\tau_{c}$ is the control input forces and moments from the propulsion system, $\tau_{disturb}^{body}$ is the external disturbance forces and moments transformed in the body frame.
The matrix components are detailed below.
$\begin{aligned} M&=\begin{bmatrix}I_{xx}&-I_{xy}&-I_{xz}&0&-mz_{g}&my_{g}\\ -I_{yx}&I_{yy}&-I_{yz}&mz_{g}&0&-mx_{g}\\ -I_{zx}&-I_{zy}&I_{zz}&-my_{g}&mx_{g}&0\\ 0&mz_{g}&-my_{g}&m&0&0\\ -mz_{g}&0&mx_{g}&0&m&0\\ my_{g}&-mx_{g}&0&0&0&m\end{bmatrix}\\ &=\begin{bmatrix}M_{11}&M_{12}\\ M_{21}&M_{22}\end{bmatrix}\end{aligned}$
where $x_{g}$, $y_{g}$, and $z_{g}$ are the distances from the origin to the center of gravity, and $I$ is the inertia dyadic about the origin coordinate.
$\begin{aligned} C(\xi)=\begin{bmatrix}-\left(M_{11}\xi_{1}+M_{12}\xi_{2}\right)^{\wedge}&-\left(M_{21}\xi_{1}+M_{22}\xi_{2}\right)^{\wedge}\\ -\left(M_{11}\xi_{1}+M_{12}\xi_{2}\right)^{\wedge}&0\end{bmatrix}\end{aligned}$
where $\xi_{1}=[p,q,r]^{\mathsf{T}}$ is the angular velocity vector, and $\xi_{2}=[u,v,w]^{\mathsf{T}}$ is the translational velocity vector. The Coriolis matrix $C(\xi)$ is skew-symmetric and represents the inertial forces in the vehicle dynamics.
@@
where $\mathrm{diag}(x)$ denotes a diagonal matrix with the elements of vector $x$ on its main diagonal. The damping terms model hydrodynamic effects such as potential damping, skin friction, wave drift damping, vortex shedding, and lifting forces. Depending on the surge velocity, either linear or nonlinear damping terms dominate. For non-coupled motion, a diagonal damping structure can be assumed, simplifying the dynamics.
Combining Eq. (5) with Euler-Poincaré equations, we arrive at the rigid body equation of motion
@@
where $F\in\mathfrak{g}^{*}$ is the external force applied to the body-fixed principal axes.
@@ III-A Error-State Convex MPC
We develop a convex error-state MPC by linearizing the error and motion dynamics in the Lie algebra, coupled with a make-up residual term from an online estimator.
Let the desired trajectory be $X_{d,t}\in\mathrm{SE}(3)$, and define the left-invariant error [barrau2016invariant] as:
@@
Define the system state $x_{t}=[\psi_{t},\xi_{t}]^{\mathsf{T}}$, leading to the linearized dynamics: The modified hydrodynamic model becomes:
$\dot{x}_{t}=A_{t}x_{t}+B_{t}\tau+h_{t}+\tilde{h}_{t},$ (15)
@@
$h_{t}$ $=\begin{bmatrix}-\xi_{d,t}\\[4.0pt] b_{t}\end{bmatrix},\quad\tilde{h}_{t}=\begin{bmatrix}\delta_{t}^{e}\\[4.0pt] \delta_{t}^{d}\end{bmatrix}$.
Here, $\delta_{t}^{e}$ represents the accumulated linearization residuals in the error-state dynamics, capturing higher-order terms neglected in the first-order approximation around $\xi_{d,t}$; $\delta_{t}^{d}$ denotes the external disturbances and unmodeled hydrodynamic effects.
Output and Cost Function
Define the tracking output:
@@
The cost function is formulated as:
@@
where $P$, $Q$, and $R$ are semi-positive definite weighting matrices.
Discrete-Time MPC Formulation
@@
where $A_{k}=I+A_{t_{k}}\Delta t$, $B_{k}=B_{t_{k}}\Delta t$, $h_{k}=h_{t_{k}}\Delta t$, and and $\tilde{h}_{k}=\tilde{h}_{t_{k}}\Delta t$. The resulting QP can be efficiently solved using the OSQP solver.
Input: Total timesteps $T$; MPC prediction horizon NN; feature map $\Phi_{f^{\star}}()$; step size $\eta$; regularizer $\lambda$
Output: Predicted residual sequence $\{\hat{\tilde{h}}_{t+k|t}\}_{k=0}^{N-1}$
1: Initialize system state $x_{1}=[\psi_{1},\xi_{1}]^{\top}$; set weight matrix $\hat{\mathbf{A}}_{1}\leftarrow 0$; set $\hat{\tilde{h}}_{0|0}\leftarrow 0$.
for $t=1$ to $T$ do
  2: Generate predicted residual sequence for MPC:
  $\hat{\tilde{h}}_{t+k|t}=\hat{\mathbf{A}}_{t}^{\top}\,\Phi_{f^{\star}}\!\left(Z_{t+k|t}\right)$,    $k=0,\dots,N-1$.
  Use $\{\hat{\tilde{h}}_{t+k|t}\}$ in the MPC problem (20) to compute $\{\tau_{t+k|t}\}$ and apply the first input $\tau_{t}$.
 3: Estimate the residual term $\tilde{h}_{t}$ from (21).
 4: Compute the buffered loss using (22).
 5: Update the weight matrix:
  $\hat{\mathbf{A}}_{t+1}\leftarrow\hat{\mathbf{A}}_{t}-\eta\,\nabla_{\hat{\mathbf{A}}_{t}}\ell_{t}$.
Algorithm 1 MPC with Fourier-based Online Learning
@@ III-B Fourier-based Online Learning
At each time step $t$, we estimate the residual term $\tilde{h}_{t}$ by comparing the true and nominal state derivatives:
$\tilde{h}_{t}=\dot{x}_{t}-(A_{t}x_{t}+B_{t}\tau_{t}+h_{t}),$ (21)
where $\dot{x}_{t}=(x_{t+1}-x_{t})/\Delta t$ is obtained from discrete-time measurements using GPS and IMU fusion[singh2025gps], and $(A_{t}x_{t}+B_{t}\tau_{t}+h_{t})$ is the nominal model from (15). The residual $\tilde{h}_{t}=[\delta_{t}^{e},\delta_{t}^{d}]^{\top}$ captures both internal linearization errors and external disturbances.
@@
$Z_{t}=(X_{t},\,t),\qquad\hat{\tilde{h}}_{t}=\hat{\mathbf{A}}^{\top}\,\Phi_{f^{\star}}(Z_{t}),$
where $\Phi_{f^{\star}}(\cdot)$ is the Fourier feature map, and weight matrix $\hat{\mathbf{A}}$ is updated online using samples stored in a sliding buffer $\mathcal{B}$. For prediction inside MPC we use $Z_{t+k|t}=(X_{t+k|t},\,t+k)$, so $\hat{\tilde{h}}_{t+k|t}=\hat{\mathbf{A}}_{t}^{\top}\Phi_{f^{\star}}(Z_{t+k|t})$.
In initialization, setting the weight matrix $\hat{\mathbf{A}}{=}0$ makes the method neutral to unknown residual patterns at deployment and avoids arbitrarily large initial predictions that can destabilize MPC.
@@
$\ell_{t}\;=\;\frac{1}{S_{t}}\!\sum_{j=t-S_{t}+1}^{t}\Big(\,\|\tilde{h}_{j}-\hat{\tilde{h}}_{j}\|^{2}+\lambda\,\|\hat{\tilde{h}}_{j}-\hat{\tilde{h}}_{j-1}\|^{2}\Big),$ (22)
where $S_{t}=\min(N,\,t)$ uses up to the latest $N$ samples. In the first $N$ steps, since fewer than $N$ samples are available, the loss in (22) automatically averages over $t$ samples. We set $\hat{\tilde{h}}_{0}\!=0$ for the smoothness term when $j=1$.
Input: Dataset $\{(Z_{t},h_{t})\}_{t=1}^{T}$; initial amplitudes $\hat{\mathbf{A}}$; initial frequency set $\{f_{j}\}_{j=1}^{M}$; feature map $\Phi_{f}()$ inner steps $n_{\text{inner}}$; epochs NN; learning rate $\eta$; regularization $\lambda$
Output: Optimized Fourier feature map $\Phi_{f^{\star}}()$
for $\text{epoch}=1$ to NN do
  Freeze $\{f_{j}\}$; unfreeze $\hat{\mathbf{A}}$
 for $k=1$ to $n_{\text{inner}}$ do
     foreach $(Z_{t},h_{t})$ in dataset do
        $\hat{h}_{t}\leftarrow\hat{\mathbf{A}}\,\Phi_{f}(Z_{t})$
        $L_{\text{inner}}\leftarrow\|\hat{h}_{t}-h_{t}\|_{1}+\lambda\|\hat{\mathbf{A}}\|_{1}$
        Update $\hat{\mathbf{A}}\leftarrow\hat{\mathbf{A}}-\eta\,\nabla_{\hat{\mathbf{A}}}L_{\text{inner}}$
 Freeze $\hat{\mathbf{A}}$; unfreeze $\{f_{j}\}$
 foreach $(Z_{t},h_{t})$ in dataset do
     $\hat{h}_{t}\leftarrow\hat{\mathbf{A}}\,\Phi_{f}(Z_{t})$
     $L_{\text{outer}}\leftarrow\|\hat{h}_{t}-h_{t}\|_{1}$
     Update $\{f_{j}\}\leftarrow\{f_{j}\}-\eta\,\nabla_{f_{j}}L_{\text{outer}}$
Perform PCA on $\{\Phi_{f}(Z_{t})\}_{t}$ and select the top-3030 features as $f^{\star}$, return the feature map $\Phi_{f^{\star}}(\cdot)$.

### v2

# Robust Trajectory Tracking of Autonomous Surface Vehicle via Lie Algebraic Online MPC

@@ Abstract
Autonomous surface vehicles (ASVs) are influenced by environmental disturbances such as wind and waves, making accurate trajectory tracking a persistent challenge in dynamic marine conditions. In this paper, we propose an efficient controller for trajectory tracking of marine vehicles under unknown disturbances by combining a convex error-state MPC on the Lie group augmented by an online learning module to compensate for these disturbances in real time. This design enables adaptive and robust tracking control while maintaining computational efficiency. Extensive evaluations in the Virtual RobotX (VRX) simulator, and real-world field experiments demonstrate that our method achieves superior tracking accuracy under various disturbance scenarios compared with existing approaches.
@@ I Introduction
The autonomous surface vehicle (ASV) has played an increasingly important role in the application of marine robots, undertaking a wide range of tasks such as hydrographic surveying, environmental monitoring, and water-quality samplings [22]. For surface vehicles, the time-varying hydrodynamic coefficients and external disturbances pose an obstacle to the controller’s trajectory-tracking ability, especially for underactuated ASV [13]. For example, winds, waves, currents, and inaccurate dynamics modeling can all disturb accurate tracking trajectory [7], which can lead to undesired consequences such as grounding, collision risks, or potential structural damage to the hull, thereby degrading reliability and operational safety.
Classical controllers like PID have been applied on marine robots [13], however, their performance is constrained under strong nonlinearities and time-varying disturbances. In contrast, Model Predictive Control (MPC) has demonstrated effectiveness in trajectory tracking, particularly in more complex environments [5] [21]. However, the non-linear hydrodynamics characteristics mean that the computation load of real-time MPC is heavy, resulting in a long response time to control signals, and a low motion frequency [16]. Moreover, small computing units lack the capacity to handle such heavy workloads, while large ones consume excessive space and energy, making them impractical for a compact ASV.
Multiple approaches have been put forward to alleviate the computational burden. In [12], MPC with projection neural network performs efficient constrained optimization problems with parallel computational capability. In [1], adaptive MPC leverages multiple approximated linear models to simplify the calculation. Another approach is to achieve geometric control [8], based on a Lie group framework, to exploit symmetry in the problem. In this concept, Lie Algebraic MPC (Lie-MPC) is proposed [19], and applied on ASV [10]. Lie-MPC reduces computation time, making online learning for non-linear hydrodynamics applicable to ASV.
@@
In parallel, several online learning methods have been proposed. In [20], a neural network–based adaptive module is incorporated into a model‐based controller to compensate for large wave forces. In [14],the Deep Neural Network (DNN) is first trained offline to learn the model dynamics, then it is updated online to fit unknown disturbances. Similarly, in [15], a DNN is used to learn the system dynamics, and an Extended State Observer (ESO) measures error and makes up for residual dynamics estimation in the kinematic control law.
However, these neural‐network approaches generally require GPU resources for offline pre‐training on large datasets, which increases computational cost. Moreover, like in [15], the online-learning pipeline is used to directly change the system dynamics, making it difficult to generalize across different system setups. In [23], all disturbances are unified into a single locally Lipschitz term within the MPC dynamics and estimated through feature-based modeling. While this approach has been demonstrated on aerial vehicles and shown to reduce computational burden, its applicability to ASVs and its interaction with a Lie-MPC controller remain unexplored.
To mitigate the research gap, the main contributions of this work are as follows:
@@
The remainder of this paper is organized as follows. Section II provides the mathematical preliminaries. Section III presents the details of our proposed method. Experiments are presented in Section IV. Finally, Section V concludes the paper and discusses future studies.
@@ II-A Dynamics on Lie Group
This section briefly introduces the rigid body dynamics on the Lie group $\mathrm{SE}(3)$. For more backgrounds on matrix Lie groups, refer to [9, 18, 3].
Let $\mathcal{G}$ be an $n$-dimensional matrix Lie group with its associated Lie algebra $\mathfrak{g}$. We define the following isomorphism
@@
To describe the 3D rigid body motion of the vehicle, we represent the rotation by matrix $R\in\mathrm{SO}(3)$ and the position vector by $p\in\mathbb{R}^{3}$ and have the full state as:
@@
$\xi=\begin{bmatrix}\omega\\ v\end{bmatrix},\quad\xi^{\wedge}=\begin{bmatrix}\omega^{\wedge}&v\\ 0&0\end{bmatrix}\in\mathfrak{se}(3).$ (4)
The adjoint map of $\mathrm{SE}(3)$ has the matrix representation:
$\mathrm{Ad}_{X}=\begin{bmatrix}R&0\\ p^{\wedge}R&R\end{bmatrix},\qquad\mathrm{ad}_{\xi}=\begin{bmatrix}\omega^{\wedge}&0\\ v^{\wedge}&\omega^{\wedge}\end{bmatrix}$, (5)
where $\mathrm{Ad}_{X}$ describes the change of frame of Lie algebra, and $\mathrm{ad}_{\xi}$ is defined by $\mathrm{ad}_{\xi}\eta=[\xi,\eta]$ with $\cdot$$\cdot$] the Lie bracket.
@@ II-B Dynamics of ASV in Fossen’s model
We will use a 6-DOF motion equations based on Fossen’s model [6] to describe the ASV, with the the prescribed Lie group’s ordering of twist vector $\xi=[\omega,v]$. The dynamics are represented in the body frame following the convention of Euler Poincare equation for single rigid body:
$M_{RB}\dot{\xi}$ $+C_{RB}(\xi)\xi+M_{AM}\dot{\xi}_{r}+C_{AM}(\xi_{r})\xi_{r}$
$+D(\xi_{r})\xi_{r}+g(\eta)=\tau_{c}+\tau_{disturb}^{body}$ (6)
with $M_{(\cdot)}$ the mass matrix, $C_{(\cdot)}$ the Coriolis matrix. The subscript $RB$ refers to the rigid-body contribution, while $AM$ means “added mass”, which represents the inertia of fluid accelerated by the vehicle motions. $D(\xi_{r})$ is the hydrodynamic damping (drag) matrix. $g(\eta)$ is the hydrostatic restoring force and moment vector, and here we assume the vehicle to be neutrally buoyant, so the hydrostatic term $g(\eta)$ becomes negligible. $\xi_{r}=\xi-\xi_{c}$ is the relative twist with respect to the ocean current $\xi_{c}$. $\tau_{c}$ is the six-dimensional control input from the propulsion system, $\tau_{disturb}^{body}$ is the six-dimensional external disturbance transformed in the body frame.
The mass matrix corresponding to $\xi$ is given by
$M=\begin{bmatrix}M_{11}&M_{12}\\ M_{21}&M_{22}\end{bmatrix},\quad M_{ij}\in\mathbb{R}^{3\times 3},$ (7)
there the submatrices are defined as $M_{11}$ is a positive definite inertia matrix in body frame, $M_{12}=M_{21}^{\top}=-m\,r_{g}^{\wedge}$, and $M_{22}=mI_{3}$, where $r_{g}=[x_{g},y_{g},z_{g}]^{\mathsf{T}}$ denotes the position vector of the center of gravity expressed in the body-fixed frame. $(\cdot)^{\wedge}:\mathbb{R}^{3}\rightarrow\mathfrak{so}(3)$ denotes the wedge operator that maps a vector to its corresponding skew-symmetric matrix. The Coriolis matrix is given by
$C(\xi)=\begin{bmatrix}-\left(M_{11}\omega+M_{12}v\right)^{\wedge}&-\left(M_{21}\omega+M_{22}v\right)^{\wedge}\\ -\left(M_{11}\omega+M_{12}v\right)^{\wedge}&0\end{bmatrix}$ (8)
which is a skew-symmetric that represents the “inertial forces” in the vehicle dynamics. The damping matrix is given by
@@
The damping terms model hydrodynamic effects such as potential damping, skin friction, wave drift damping, vortex shedding, and lifting forces. Depending on the surge velocity, either linear or nonlinear damping terms dominate. For non-coupled motion, a diagonal damping structure can be assumed, simplifying the dynamics.
Combining (5) with (6), have the full dynamics
@@
with $F\in\mathfrak{se}^{*}(3)$ the external force in the body frame.
@@ III-A Error-State Convex MPC
We develop a convex error-state MPC by linearizing the tracking error dynamics in the Lie algebra and learn the residual dynamics to compensate for the modeling error.
Let the desired trajectory be $X_{d,t}\in\mathrm{SE}(3)$, and define the left-invariant error [2] on matrix Lie group as:
@@
Define the system state $x_{t}=[\psi_{t},\xi_{t}]^{\mathsf{T}}$, and the linearized hydrodynamic model becomes:
$\dot{x}_{t}=A_{t}x_{t}+B_{t}\tau+h_{t}+\delta_{t},$ (15)
with each component defined by:
@@
$h_{t}$ $=\begin{bmatrix}-\xi_{d,t}\\[4.0pt] b_{t}\end{bmatrix},\quad\delta_{t}=\begin{bmatrix}\delta_{t}^{e}\\[4.0pt] \delta_{t}^{d}\end{bmatrix}$.
Here, $\delta_{t}^{e}$ represents the higher-order linearization residuals in the error-state dynamics, capturing higher-order terms neglected in the first-order approximation around $\xi_{d,t}$, $\delta_{t}^{d}$ denotes the external disturbances and unmodeled hydrodynamic effects.
For cost function, we first define the tracking output:
@@
Then the cost function is formulated as:
@@
with $P$, $Q$, and $R$ the positive definite weighting matrices.
@@
where $A_{k}=I+A_{t_{k}}\Delta t$, $B_{k}=B_{t_{k}}\Delta t$, $h_{k}=h_{t_{k}}\Delta t$, and and $\delta_{k}=\delta_{t_{k}}\Delta t$. The resulting QP can be efficiently solved using the OSQP solver.
Algorithm 1 Fourier Online MPC
1:Total timesteps $T$; MPC prediction horizon NN; feature map $\Phi(\cdot)$; step size $\eta$; regularizer $\lambda$
[…]
5:for $t=1,\ldots,T$ do
6:  $\{\hat{\delta}_{t+k|t}\}_{k=1}^{H}\;\leftarrow\;\{\hat{C}_{t}^{\top}\Phi(Z_{t+k|t})\}_{k=1}^{H}\triangleright$ Predict Residual
7:  $\tau_{t}\;\leftarrow\;\mathrm{MPC}\!\big(x_{t},\,\{\hat{\delta}_{t+k}\}_{k=1}^{H}\big)$ $\triangleright$ Solve (20)
8:  $\delta_{t}=\dot{x}_{t}-(A_{t}x_{t}+B_{t}\tau_{t}+h_{t})$ $\triangleright$ Estimate residual
[…]
10:  $\hat{C}_{t+1}\leftarrow\hat{C}_{t}-\eta\nabla_{\hat{C}_{t}}\ell_{t}$ $\triangleright$ Update weights
11:end for
@@ III-B Fourier-based Online Learning
At each time step $t$, we estimate the residual term $\delta_{t}$ by comparing the measured and nominal state derivatives:
$\delta_{t}=\dot{x}_{t}-(A_{t}x_{t}+B_{t}\tau_{t}+h_{t}),$ (21)
where $\dot{x}_{t}=(x_{t+1}-x_{t})/\Delta t$ is obtained from discrete-time measurements using GPS and IMU fusion[17], and $(A_{t}x_{t}+B_{t}\tau_{t}+h_{t})$ is the nominal model from (15). The residual $\delta_{t}=[\delta_{t}^{e},\delta_{t}^{d}]^{\top}$ captures both internal linearization errors and external disturbances.
@@
$Z_{t}=(x_{t},t),\qquad\hat{\delta}_{t}={C}^{\top}\,\Phi(Z_{t}),$ (22)
where $\Phi(\cdot):\mathbb{R}^{12}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$ is the Fourier feature map that maps the time-dependent robot state to a $d$-dimensional feature. The coefficient matrix C{C} is updated online using samples stored in a sliding buffer $\mathcal{B}$.
@@
$\ell_{t}\;=\;\frac{1}{N}\!\sum_{j=t-N+1}^{t}\Big(\,\|\delta_{j}-\hat{\delta}_{j}\|^{2}+\lambda\,\|\hat{\delta}_{j}-\hat{\delta}_{j-1}\|^{2}\Big),$ (23)
We set $\hat{\delta}_{0}\!=0$ for the smoothness term when $j=1$. The optimization problem is formulated as
$\hat{C}=\arg\min_{C}\quad\ell_{t}$ (24)
III-C Offline Bi-level Feature Extraction
Inspired by the Fourier-based nature of wave and wind disturbances [11], the features we use are all Fourier-based. However, directly using a large number of random Fourier features for disturbance estimation causes many of them fail to capture the dominant frequency components and spatial correlations of the residual term. To address this issue, we perform feature extraction to identify and retain components that contribute significantly to the residual representation. This reduces both the computational cost and the risk of overfitting while preserving model accuracy.
$\Phi(Z_{t})=\big[\sin(f_{j}\,z_{t,i}),\;\cos(f_{j}\,z_{t,i})\big]_{i=1,\dots,d;\,j=1,\dots,m}$ (25)